If (2x 2 -3x+1)(2x 2 +5x+1) =9x 2 , then eqn has a) Four real roots b) Two real and two imaginary c) Four imaginary roots d) none


If  (2x2-3x+1)(2x2+5x+1) =9x2, then eqn has


a)     Four real roots


b)    Two real and two imaginary


c)     Four imaginary roots


d)    none


Grade:12

1 Answers

jitender lakhanpal
12 years ago

Dear pooja, 

correct option is a)

solution

 

(2x2-3x+1)(2x2+5x+1) =9x2

divide the equation by x^2 where x not equal to 0

(2x-3+1/x)(2x+5+1/x) = 9

 let  2x+1/x = y  put in the equation 

(y-3)(y+5) = 9

y2+2 y - 24 = 0 

solving the equation we get  y = -6 and y = 4 

so we get   

 2x+1/x = -6        and 

 2x+1/x = 4      we get 

2x2+6x+1=0    finding discriminant we get it as  +ve so 2 real roots  

2x2-4x+1=0 and again finding discriminant we get it as +ve so 2 real roots   so the total of 4 real roots 

 

 

 

 

 

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  

jitender

askiitians expert